OpenStudy (anonymous):

could someone help me in how to solve partial fractions Integrate(4w^2+3w^2+9/w^8-9w)dw

OpenStudy (anonymous):

First factor the denominator.

OpenStudy (anonymous):

ok then i'm sorry could please explain it to with some more detail thanks

OpenStudy (anonymous):

i really don't know how to do this for me its like from scratch

OpenStudy (anonymous):

\[ \frac{4w^2+3w^2+9}{w^8-9w}=\frac{7w^2+9}{w(w^7-9)} \]

OpenStudy (anonymous):

Each denominator gets its own fraction: \[ \frac{7w^2+9}{w(w^7-9)} = \frac{A}{w}+\frac{B}{w^7-9} \]

OpenStudy (anonymous):

how did you got 7w^2+9 for nominator?

OpenStudy (anonymous):

Then you try to add the fractions as are: \[ \frac{7w^2+9}{w(w^7-9)} = \frac{A}{w}+\frac{B}{w^7-9}=\frac{A(w^7-9)+Bw}{w(w^7-9)} \]

OpenStudy (anonymous):

You wrote:\[ 4w^2+3w^2+9 \]Add like terms: \[ 7w^2+9 \]

OpenStudy (anonymous):

i'm sorry i made a mistake but 3w not 3w^2

OpenStudy (anonymous):

but i got it the point

OpenStudy (anonymous):

In this case fraction decomposition won't work because there is no solution for \(A\) and \(B\).

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Unless you wrote it incorrectly...

OpenStudy (anonymous):

no i wrote it correctly but can't we solve for A and B?

OpenStudy (anonymous):

Doesn't look like it: \[ 4w^2+2w+9=Aw^7+Bw-9A \]

OpenStudy (anonymous):

yeah i guess it wouldn't work like this

OpenStudy (anonymous):

Shouldn't it be \(\color{blue}{\text{Originally Posted by}}\) @wio \[\frac{7w^2+9}{w(w^7-9)} = \frac{A}{w}+\frac{\color{red}{Bw^6+Cw^5+Dw^4+Ew^3+Fw^2+Gw+H}}{w^7-9}~~? \]\(\color{blue}{\text{End of Quote}}\)

OpenStudy (anonymous):

A lot of those constants will likely disappear, I think.

OpenStudy (anonymous):

The \(Bx+C\) comes when you have denominators with multiplicity greater than 1.

OpenStudy (anonymous):

Right, but if you have a quadratic factor in the denominator, the numerator of the partial fraction with that denominator will be a linear factor.

OpenStudy (anonymous):

So in general, if you have an \(n\)-th degree factor in the denominator, you should have an \(n-1\)-th degree factor in the numerator. In the case of greater-than-1 multiplicity, you can use \(Bx+C\) for a factor with mult. 2, but you'll find that \(B=0\).

OpenStudy (anonymous):

\[\begin{align*}\frac{4w^2+3w+9}{w(w^7-9)} &= \frac{A}{w}+\frac{Bw^6+Cw^5+Dw^4+Ew^3+Fw^2+Gw+H}{w^7-9}\\ 4w^2+3w+9&=Aw^7-9A+Bw^7+Cw^6+Dw^5+Ew^4+Fw^3+Gw^2+Hw\\ &=(A+B)w^7+Cw^6+Dw^5+Ew^4+Fw^3+Gw^2+Hw-9A \end{align*}\] which gives the system \[\begin{cases} A+B=0\\C=0\\D=0\\E=0\\F=0\\G=4\\H=3\\-9A=9 \end{cases}~~\Rightarrow~~A=-1,~B=1\]